# Euclid preparation

### v1

@@ 1 Introduction
In this paper we present the simulated spectroscopic skies that have been produced by the Euclid Collaboration to face these challenges. These are based on the largest set of simulated dark matter haloes on the past light cone ever produced. We discuss two sets of simulated Euclid spectroscopic skies (whose properties are reported in Table 1), performed with the Pinocchio code (Monaco et al. 2002) for a flat $\Lambda$ cold dark matter ($\Lambda$CDM) model compatible with Planck results (Planck Collaboration: Aghanim et al. 2020). The first set of 3500 realisations, named Geppetto, is based on a relatively small box of $1.2\,h^{-1}\,\text{Gpc}$, sampled with $2160^{3}$ particles to achieve a minimum halo mass of $1.5\times 10^{11}\,h^{-1}\,M_{\odot}$. Each light-cone covers a circle of 30 degrees of radius on the sky, covering 2763 ${\rm deg}^{2}$, a bit larger than the area planned to be covered by Euclid’s DR1 (Euclid Collaboration: Scaramella et al. 2022), and extends to redshift $z=2$. The second set of 1000 realisations (named EuclidLargeBox) is based on a $3.38\,h^{-1}\,\text{Gpc}$ box, sampled by $6144^{3}$ particles and achieving the same mass resolution as Geppetto. The light-cones cover half of the sky, excluding a zone of avoidance of the Milky Way, for a semi-aperture angle of $70$ deg, and extend to $z=4$.
@@ 2 Context
Our starting point is the upcoming spectroscopic sample of Euclid (Euclid Collaboration: Mellier et al. 2025), consisting of $\sim$ 30 million galaxies with reliable redshift measurement in the range $1.80.9<z<$ 1.8 over 14 000 ${\rm deg}^{2}$. Systematic effects for the measurement of the galaxy density are represented by constructing a ‘visibility mask’, a function of sky position and redshift that quantifies, to the best of our knowledge, the probability that a galaxy gets a reliable redshift measurement given its properties and the local observing conditions. This visibility mask is applied to a dense set of unclustered galaxies to obtain a ‘random catalogue’, that is used as a reference to compute the galaxy density contrast, thus mitigating systematic effects induced by a non-uniform selection probability. All estimators of summary statistics that are based on the galaxy density rely on this random catalogue.
@@
A numerical covariance matrix is affected by noise as it is sampled by a finite number of mocks; for Euclid, the requirement on its accuracy is that parameter uncertainties (for the final sample after the nominal six years of the mission) should vary by less than 10%, with respect to an infinite number of mocks. In Euclid Collaboration: Sanchez et al. (in prep.), we estimate that 3500 mocks are needed to have an accurate, brute-force numerical covariance matrix. This number is mostly determined by the total number of measurements for which the covariance is computed and by the accuracy requirement. That paper also presents tests of several methods to de-noise a numerical covariance matrix, demonstrating that $\sim$ 100 mocks may be sufficient to achieve good accuracy (see also Fumagalli et al. 2022) when addressing the relatively limited data vector of 2-point correlation functions (in configuration or Fourier space). However, this figure may be optimistic: the inclusion of 3-point statistics will further increase the length of the data vector, moreover a precise characterisation of systematic effects requires to analyse a large number of mock catalogues (e.g. Euclid Collaboration: Risso et al., in prep; Euclid Collaboration: Lee et al., in prep.). Under these premises, we consider 1000 as a minimal requirement on the number of realisations for a set of simulated catalogues for Euclid.
[…]
@@ 4 Dark matter catalogues and light-cones
As discussed above, we produced two sets of simulations aimed at achieving a large number of realizations based on a small box (Geppetto) and a large box that can contain a survey, with a more limited number of realizations (EuclidLargeBox). The main properties of these sets are reported in Table 1. Both sets are based on a $\Lambda$CDM cosmology with parameters close to Planck Collaboration: Aghanim et al. (2020): $\Omega_{\rm m}=0.32$, $\Omega_{\Lambda}=0.68$, $\Omega_{\rm b}=0.049$, $h=0.67$, $\sigma_{8}=0.83$, $n_{\rm s}=0.96$. These parameters are consistent with Flagship 2 simulation and mock galaxy catalogue; in the Flagship 2 simulation we have $\Omega_{\rm m}=0.319$ and massive neutrinos, but these differences have no practical effect for the present purposes. For each simulation we produced a few dark matter halo catalogues in periodic boxes and a complete catalogue in the past light cone, with a sky footprint consisting of a circle with a given radius. Moreover, merger trees are available for all haloes. The minimum halo mass is set to 10 particles; this would be an unacceptably low value for an $N$-body simulation, but it is acceptable for a semi-analytic code where haloes are related to peaks of the inverse collapse time and each halo starts with one particle.
Table 1: Main properties of the two sets of mock catalogues presented in this paper, Geppetto and EuclidLargeBox, together with the Minerva-like set used in Oddo et al. (2020) and NewClusterMocks used in Fumagalli et al. (2021). Columns give the simulation box side $L_{\rm s}$, the number of particles in the simulation $N_{\rm part}$, the number of realisations $N_{\rm real}$, the minimum halo mass $M_{\rm h}$, the simulation volume $V_{\rm s}$ and its total over the realisations, the semi-aperture of the light-cone $\theta$, the survey area, the starting redshift of the light-cone $z_{\rm start}$, the volume of a single light-cone $V_{\rm lc}$ and its total over the realisations.
@@ 4.1 Geppetto simulations
The first set of $N_{\rm real}=3500$ simulations, called Geppetto, was designed to be massive but relatively inexpensive, to produce a brute-force numerical covariance for the target sample of $\mathrm{H}\,\alpha$ emitters. It is based on a relatively small box of side $L_{\rm s}=1200$$h^{-1}\,\text{Mpc}$, sampled with $N_{\rm part}=2160^{3}$ particles. With the given cosmology, the particle mass is $1.52\times 10^{10}$$h^{-1}\,M_{\odot}$, and the smallest halo has a mass of 10 particles or $M_{\rm h}=1.52\times 10^{11}$$h^{-1}\,M_{\odot}$. The light-cones start at $z_{\rm start}=2$ and cover a circle of radius of $\theta=30^{\circ}$, for a total area of $2\pi\,(1-\cos\theta)$ steradians, equal to 2763 ${\rm deg}^{2}$.333 Noticeably, this specific sky footprint implies that the comoving volume of the light-cone is a 3D cone itself; however, the word ‘cone’ in the two cases refers to a spacetime and to a volume in 3D space, thus assuming a pretty different meaning. Outputs at fixed time are available at redshifts 2, 1.8, 1.5, 1.35, 1, 0.9, 0.5, and 0.
@@ 4.2 The problem of replications
To highlight this effect, we computed the power spectrum of the dark matter haloes of the Geppetto light-cones, with masses $>10^{12}$$h^{-1}\,M_{\odot}$. In order to correct for the small difference in halo mass function between the first 600 light-cones and the remaining ones (Fig 2), we set the lower limits in halo mass at $10^{12}$$h^{-1}\,M_{\odot}$, corresponding to 66 particles, in the first 600 runs and to $0.99\times 10^{12}$$h^{-1}\,M_{\odot}$, corresponding to 65 particles, in the other runs. Although the difference corresponds to only 1 particle, it is sufficient to ensure a consistent amplitude of the halo power spectrum. The power spectrum was computed using the Euclid ground-segment code (Euclid Collaboration: Salvalaggio et al., in prep.) that implements a Yamamoto–Bianchi estimator (Yamamoto et al. 2006; Bianchi et al. 2015). We produced catalogues of DM haloes in redshift space, in the redshift range $z\in[0.9,1.1]$, and immersed them in boxes of side $L_{\rm m}=7.2$$h^{-1}\,\text{Gpc}$. The resulting power spectrum is shown in Fig. 3, where we show averages for 100 mocks performed with V4 and with cone axis aligned with the main diagonal (orange lines) and 100 mocks performed with V5 with random direction of the cone axis (blue lines) to mitigate the replication problem. The figure also reports the fundamental modes of the simulation box $k_{\rm fs}$ and of the measurement box $k_{\rm fm}$. The beating due to the oversampling of the Fourier space is very evident especially in the orange lines, and the randomisation of the cone direction mitigates but does not remove the oscillations (averaging over more realisations further dampens these oscillations without fully removing them).
@@
4.2.1 EuclidLargeBox simulations
[…]
To handle the memory overhead of the fragmentation and its variations over many realisations, it was necessary to allocate 250 bytes per particle, for a total memory requirement of 52 TB. After a few experiments, we decided to run the code on 24 fat nodes of the machine Galileo100@CINECA, equipped with 3 TB each and 48 cores per node. Each run took on average of 3 hours 45 minutes, for a computational cost of 4276 core-h per run, amounting to a total of about 4.3 million core-h.
@@ 5 From haloes to galaxies
$N_{\rm cen}(M_{\rm h},z\,|\,f_{{\rm H}\alpha}\geq f_{\rm lim})$ $=$ #$\frac{\rm\#\ of\ central\ galaxies}{\rm\#\ of\ haloes}\,,$ (1)
$N_{\rm sat}(M_{\rm h},z\,|\,f_{{\rm H}\alpha}\geq f_{\rm lim})$ $=$ #$\frac{\rm\#\ of\ satellite\ galaxies}{\rm\#\ of\ haloes}\,,$ (2)
@@
[…]
@@
Pinocchio and $N$-body halo masses are not expected to be equivalent, both because Pinocchio’s claimed accuracy is $\sim$ 5% in the halo mass function and because it has been calibrated to reproduce haloes found with the friends-of-friends algorithm, not those found with the Rockstar halo finder (Behroozi et al. 2013) used in the Flagship simulation (see Euclid Collaboration: Castro et al. 2023 for a discussion of the differences among halo finders). Moreover, Pinocchio is able to reproduce the linear halo bias to within a few percent (Paranjape et al. 2013; Munari et al. 2017b), and this difference in clustering amplitude can be absorbed by halo calibration. This was done in Oddo et al. (2020), where an optimal reproduction of the halo bispectrum and its covariance was achieved by calibrating the mass cuts of simulated and Pinocchio haloes so as to have the same amplitude of the power spectrum. To apply this HOD to our halo light-cones, we calibrate halo masses with a clustering matching (CM) procedure, described in detail in the Appendix, and different from the abundance matching (AM) procedure that would guarantee consistency in the galaxy number density. Once the clustering level is matched, matching of number density is achieved by multiplying the HOD curves $N_{\rm cen}$ and $N_{\rm sat}$ by the ratio of a target number density and the measured one. Indeed, as long as the largest value of $N_{\rm cen}$ remains below unity, these curves can be shifted up or down with no effect on the clustering amplitude (apart the obviously different shot noise level), as the changing amplitude leads to a different sampling of the parent halo distribution, keeping the ratios of haloes of different mass constant.
@@
The number density of the Flagship catalogue was calibrated to reproduce model 3 of Pozzetti et al. (2016), achieving an accuracy of roughly 5%, larger than the expected sample variance though not larger than the observational uncertainty. As a target number density we then used that of the Flagship catalogue. We fitted the ratio of the number density obtained with CM, averaged over 100 mocks, and the Flagship density with a 4th-order polynomial, and used this fitted ratio to modulate the two HOD curves as a function of redshift. We give the resulting number density $n_{\rm g}$ in Fig. 6, compared with that of the Flagship mock on the same $30^{\circ}$ radius footprint and with model 3, for three flux limits, $f_{{\rm H}\alpha}>0.5f_{0}$, $f_{0}$, and $1.5f_{0}$, amounting to $1$, $2$ and $3\times 10^{-16}$ erg s-1 cm-2. For the Geppetto and EuclidLargeMocks sets we report the standard deviation of the number density as a shaded area; the same standard deviation (from the EuclidLargeMocks) is assigned to the Flagship and shown as an errorbar. The three lower panels report the difference of the number densities with respect to model 3 predictions (not reported in the above panel) for the three flux limits. The Flagship mock fluctuates around the Geppetto and EuclidLargeMocks average values in a way that is compatible with the expected standard deviation, with some difference at the lowest redshift; at low flux the little blips in the Flagship number density, at $z=1$ and $1.5$, is due to the issue connected to the interpolation of internal extinction mentioned in Sect. 5.

### v3

@@ 1 Introduction
[…]
@@ 2 Context
Our starting point is the upcoming spectroscopic sample of Euclid (Euclid Collaboration: Mellier et al. 2025), consisting of $\sim$ 30 million galaxies with reliable redshift measurement in the range $0.9<z<1.8$ over 14 000 ${\rm deg}^{2}$. Systematic effects for the measurement of the galaxy density are represented by constructing a ‘visibility mask’, a function of sky position and redshift that quantifies, to the best of our knowledge, the probability that a galaxy gets a reliable redshift measurement given its properties and the local observing conditions. This visibility mask is applied to a dense set of unclustered galaxies to obtain a ‘random catalogue’, that is used as a reference to compute the galaxy density contrast, thus mitigating systematic effects induced by a non-uniform selection probability. All estimators of summary statistics that are based on the galaxy density rely on this random catalogue.
@@
A numerical covariance matrix is affected by noise as it is sampled by a finite number of mocks; for Euclid, the requirement on its accuracy is that parameter uncertainties (for the final sample after the nominal six years of the mission) should vary by less than 10%, with respect to an infinite number of mocks. In Euclid Collaboration: Sanchez et al. (in prep.), we estimate that 3500 mocks are needed to have an accurate, brute-force numerical covariance matrix. This number is mostly determined by the total number of measurements for which the covariance is computed and by the accuracy requirement. That paper also presents tests of several methods to de-noise a numerical covariance matrix, demonstrating that $\sim$ 100 mocks may be sufficient to achieve good accuracy (see also Fumagalli et al. 2022) when addressing the relatively limited data vector of 2-point correlation functions (in configuration or Fourier space). However, this figure may be optimistic: the inclusion of 3-point statistics will further increase the length of the data vector, moreover a precise characterisation of systematic effects requires to analyse a large number of mock catalogues (e.g. Euclid Collaboration: Risso et al. 2025; Euclid Collaboration: Lee et al., in prep.). Under these premises, we consider 1000 as a minimal requirement on the number of realisations for a set of simulated catalogues for Euclid.
A single simulation should ideally cover a volume that can include the largest connected patch of the surveyed volume, so as to properly represent super-sample covariance; this pushes requirements on the box size beyond $3\,h^{-1}\,\text{Gpc}$. The simulation should also resolve the smallest halo that contains an observed galaxy, that (as argued in Sect. 5) is expected to be $\sim$$2\times 10^{11}$$h^{-1}\,M_{\odot}$ for the spectroscopic sample; this pushes requirements on particle mass towards $10^{9}$$h^{-1}\,M_{\odot}$. One such simulation was used to create the Flagship galaxy mock catalogue (Euclid Collaboration: Castander et al. 2025), a comprehensive catalogue representing galaxies in all aspects that are of interest for Euclid, including weak lensing, galaxy clusters, and galaxy evolution. The single Flagship numerical simulation used to create the galaxy mock catalogue costed $\sim$ 1 000 000 node-hours (corresponding to 68 000 000 core hours) on 4000 nodes of Piz Daint supercomputer, and required storage of order of petabytes. Running thousands of such simulations is out of the question.
@@ 4 Dark matter catalogues and light-cones
[…]
Table 1: Main properties of the mock catalogues. [^4] We report here Geppetto and EuclidLargeBox, together with the Minerva-like set used in Oddo et al. (2020) and NewClusterMocks used in Fumagalli et al. (2021). Columns give the simulation box side $L_{\rm s}$, the number of particles in the simulation $N_{\rm part}$, the number of realisations $N_{\rm real}$, the minimum halo mass $M_{\rm h}$, the simulation volume $V_{\rm s}$ and its total over the realisations, the semi-aperture of the light-cone $\theta$, the survey area, the starting redshift of the light-cone $z_{\rm start}$, the volume of a single light-cone $V_{\rm lc}$ and its total over the realisations.
@@ 4.1 Geppetto simulations
The first set of $N_{\rm real}=3500$ simulations, called Geppetto, was designed to be massive but relatively inexpensive, to produce a brute-force numerical covariance for the target sample of $\mathrm{H}\,\alpha$ emitters. It is based on a relatively small box of side $L_{\rm s}=1200$$h^{-1}\,\text{Mpc}$, sampled with $N_{\rm part}=2160^{3}$ particles. With the given cosmology, the particle mass is $1.52\times 10^{10}$$h^{-1}\,M_{\odot}$, and the smallest halo has a mass of 10 particles or $M_{\rm h}=1.52\times 10^{11}$$h^{-1}\,M_{\odot}$. The light-cones start at $z_{\rm start}=2$ and cover a circle of radius of $\theta=30^{\circ}$, for a total area of $2\pi\,(1-\cos\theta)$ steradians, equal to 2763 ${\rm deg}^{2}$. [^5] Noticeably, this specific sky footprint implies that the comoving volume of the light-cone is a 3D cone itself; however, the word ‘cone’ in the two cases refers to a spacetime and to a volume in 3D space, thus assuming a pretty different meaning. Outputs at fixed time are available at redshifts 2, 1.8, 1.5, 1.35, 1, 0.9, 0.5, and 0.
@@ 4.2 The problem of replications
To highlight this effect, we computed the power spectrum of the dark matter haloes of the Geppetto light-cones, with masses $10^{12}$$h^{-1}\,M_{\odot}$. In order to correct for the small difference in halo mass function between the first 600 light-cones and the remaining ones (Fig 2), we set the lower limits in halo mass at $10^{12}$$h^{-1}\,M_{\odot}$, corresponding to 66 particles, in the first 600 runs and to $0.99\times 10^{12}$$h^{-1}\,M_{\odot}$, corresponding to 65 particles, in the other runs. Although the difference corresponds to only 1 particle, it is sufficient to ensure a consistent amplitude of the halo power spectrum. The power spectrum was computed using the Euclid ground-segment code (Euclid Collaboration: Salvalaggio et al., in prep.) that implements a Yamamoto–Bianchi estimator (Yamamoto et al. 2006; Bianchi et al. 2015). We produced catalogues of DM haloes in redshift space, in the redshift range $z\in[0.9,1.1]$, and immersed them in boxes of side $L_{\rm m}=7.2$$h^{-1}\,\text{Gpc}$. The resulting power spectrum is shown in Fig. 3, where we show averages for 100 mocks performed with V4 and with cone axis aligned with the main diagonal (orange lines) and 100 mocks performed with V5 with random direction of the cone axis (blue lines) to mitigate the replication problem. The figure also reports the fundamental modes of the simulation box $k_{\rm fs}$ and of the measurement box $k_{\rm fm}$. The beating due to the oversampling of the Fourier space is very evident especially in the orange lines, and the randomisation of the cone direction mitigates but does not remove the oscillations (averaging over more realisations further dampens these oscillations without fully removing them).
@@
4.3 EuclidLargeBox simulations
The second set of $N_{\rm real}=1000$ simulations, called EuclidLargeBox, was designed to have a more limited set of much larger and more expensive simulations, not affected by the replications problem. They are based on a $L_{\rm s}=3.38$$h^{-1}\,\text{Gpc}$ box sampled with $N_{\rm part}=6144^{3}$ particles. With these choices, the particle mass is $1.48\times 10^{10}$$h^{-1}\,M_{\odot}$, very similar to that of Geppetto simulations, and the smallest output halo (with at least 10 particles) has a mass of $M_{\rm h}=1.48\times 10^{11}$$h^{-1}\,M_{\odot}$. This box is marginally smaller than that (3.6 $h^{-1}\,\text{Gpc}$) of the Flagship simulation, that is however sampled with $16\,000^{3}$ particles, reaching a particle mass of $10^{9}$$h^{-1}\,M_{\odot}$. The halo mass functions of the EuclidLargeBox set, shown in Fig. 2, are consistent to within 1% with the ones from the Geppetto mocks. These simulations produced much larger light-cones, covering half of the sky and starting at redshift $z=4$. Because the Euclid survey will never go deep in the zone of avoidance of the Milky Way, sampling Galactic latitudes always with $|b|>20$, light-cones were produced with a sky footprint of a circle with radius of $\theta=70^{\circ}$, thus reducing storage needs by 34% with respect to a really half-sky output. With the same aim of reducing storage, output in periodic boxes was limited to redshifts 0 and 1. This produced 215 GB per run, 210 TB in total, of which 80 GB per run were taken by the light-cones.
To handle the memory overhead of the fragmentation and its variations over many realisations, it was necessary to allocate 250 bytes per particle, for a total memory requirement of 52 TB. After a few experiments, we decided to run the code on 24 fat nodes of the machine Galileo100 at CINECA, equipped with 3 TB each and 48 cores per node. Each run took on average of 3 hours 45 minutes, for a computational cost of 4276 core-h per run, amounting to a total of about 4.3 million core-h.
@@ 5 From haloes to galaxies
$N_{\rm cen}(M_{\rm h},z\,|\,f_{{\rm H}\alpha}\geq f_{\rm lim})$ $=$ $\frac{\rm\#\ of\ central\ galaxies}{\rm\#\ of\ haloes}\,$, (1)
$N_{\rm sat}(M_{\rm h},z\,|\,f_{{\rm H}\alpha}\geq f_{\rm lim})$ $=$ $\frac{\rm\#\ of\ satellite\ galaxies}{\rm\#\ of\ haloes}\,$, (2)
@@
Figure 6: Comoving number density $n_{\rm g}$, in $h^{3}\,\mathrm{Mpc}^{-3}$, of galaxies as a function of redshift for Geppetto (orange), and EuclidLargeMocks (red) catalogues, compared with the Flagship catalogue (dark blue points). For the sets of Pinocchio catalogues we give the average number density as a continuous line and its standard deviation as a shaded area, while Flagship values, reported as circles, are assigned the EuclidLargeMocks variance, reported as an errorbar. We show in the top panel results for three limiting fluxes, $f_{{\rm H}\alpha}>0.5f_{0}$, $f_{0}$, and $1.5f_{0}$. The other panels give, for each limiting flux, the residuals ($n_{\rm g}/n_{\rm model\,3}-1$) in percent of simulated number densities with respect to model 3 from Pozzetti et al. (2016). The black lines denote the zero value of the residuals.
@@
Pinocchio and $N$-body halo masses are not expected to be equivalent, both because Pinocchio’s claimed accuracy is $\sim$ 5% in the halo mass function and because it has been calibrated to reproduce haloes found with the friends-of-friends algorithm, not those found with the Rockstar halo finder (Behroozi et al. 2013) used in the Flagship simulation (see Euclid Collaboration: Castro et al. 2023 for a discussion of the differences among halo finders). Moreover, Pinocchio is able to reproduce the linear halo bias to within a few percent (Paranjape et al. 2013; Munari et al. 2017b), and this difference in clustering amplitude can be absorbed by halo calibration. This was done in Oddo et al. (2020), where an optimal reproduction of the halo bispectrum and its covariance was achieved by calibrating the mass cuts of simulated and Pinocchio haloes so as to have the same amplitude of the power spectrum. To apply this HOD to our halo light-cones, we calibrate halo masses with a clustering matching (CM) procedure, described in detail in the Appendix, and different from the abundance matching (AM) procedure that would guarantee consistency in the galaxy number density. As discussed in the Appendix, this match (obtained with a simple procedure) guarantees the same clustering amplitude for the halos containing the bulk of the $\mathrm{H}\,\alpha$ galaxies and is adequate for the present purposes, but it should be made more sophisticated to better account for the mass dependence of the accuracy of halo bias. Once the clustering level is matched, matching of number density is achieved by multiplying the HOD curves $N_{\rm cen}$ and $N_{\rm sat}$ by the ratio of a target number density and the measured one. Indeed, as long as the largest value of $N_{\rm cen}$ remains below unity, these curves can be shifted up or down with no effect on the clustering amplitude (apart the obviously different shot noise level), as the changing amplitude leads to a different sampling of the parent halo distribution, keeping the ratios of haloes of different mass constant.
@@
The number density of the Flagship catalogue was calibrated to reproduce model 3 of Pozzetti et al. (2016), achieving an accuracy of roughly 5%, larger than the expected sample variance though not larger than the observational uncertainty. As a target number density we then used that of the Flagship catalogue. We fitted the ratio of the number density obtained with CM, averaged over 100 mocks, and the Flagship density with a 4th-order polynomial, and used this fitted ratio to modulate the two HOD curves as a function of redshift. We give the resulting number density $n_{\rm g}$ in Fig. 6, compared with that of the Flagship mock on the same $30^{\circ}$ radius footprint and with model 3, for three flux limits, $f_{{\rm H}\alpha}>0.5f_{0}$, $f_{0}$, and $1.5f_{0}$, amounting to $1$, 22 and $3\times 10^{-16}$ erg s-1 cm-2. For the Geppetto and EuclidLargeMocks sets we report the standard deviation of the number density as a shaded area; the same standard deviation (from the EuclidLargeMocks) is assigned to the Flagship and shown as an errorbar. The three lower panels report the difference of the number densities with respect to model 3 predictions (not reported in the above panel) for the three flux limits. The Flagship mock fluctuates around the Geppetto and EuclidLargeMocks average values in a way that is compatible with the expected standard deviation, with some difference at the lowest redshift; at low flux the little blips in the Flagship number density, at $z=1$ and $1.5$, is due to the issue connected to the interpolation of internal extinction mentioned in Sect. 5.